# Bridging Domain Gap for Flight-Ready Spaceborne Vision

### v1

@@ 5.1 Datasets
Following SPEC2021 [17], this section performs experiments using the SPEED+ dataset [15, 113] which consists of 59,960 synthetic images split in 80:20 ratio into the training and validation sets. It further includes two HIL image domains acquired from the TRON facility and a mockup model of the Tango spacecraft: 6,740 lightbox and 2,791 sunlamp images which are reserved solely for testing. In this section, all NN models are trained on the synthetic training set and evaluated on the synthetic validation set, lightbox and sunlamp test sets, and 25 labeled flight images from acquired in-space during RPO of the PRISMA mission (prisma25). Recall Fig. 1 for visualization of a few HIL domain images.
@@ 5.2 Implementations
Given the ongoing interest in Artificial Intelligence (AI) for space applications, this work assumes that hardware support for edge computing to run NN onboard satellites will increase in the upcoming years (see section 3.1.1). Therefore, this work adopts an NVIDIA Jetson Nano 4GB which contains a Quad-Core Arm® Cortex® A57 CPU and a 128-core NVIDIA Maxwell™ architecture GPU [23]. Nano is the most restrictive version of the commercially available Jetson family, which suits the limited processing power of onboard GPUs as evidenced in the NASA small spacecraft technology report [83]. Given its deprecated GPU architecture and CUDA language support, all NN training and inference on Jetson Nano are performed with full 32-bit Floating Point (FP32) precision using the C++ API of PyTorch 1.10.0.
5.2.1 Default Configuration
The default training configuration consists of the heatmap prediction head from the P2 level feature output and all data augmentations including style augmentation, DeepAugment and RandConv. The ground-truth heatmaps are Gaussian blobs with the standard deviation $\sigma=1$ pixel around the 2D keypoint locations. The NN is trained with the pixel-wise Mean Squared Error (MSE) loss defined for the predicted heatmap $\hat{\bm{\mathcal{H}}}$ and the ground-truth $\bm{\mathcal{H}}$ as
@@ 5.2.2 Metrics
Also note that for the SPEED+ HIL domains, the rotation and translation errors account for the calibration accuracy of the TRON facility. Specifically, for individual samples, the HIL errors are defined as [17]
@@
Note that in the overall pose estimation scenario (section 4.1), the information on the relative position of the target is mostly included in the RoI detected from either the navigation filter or a separate object detection NN. Since the ground-truth RoIs are used for image pre-processing in this section, mean translation errors are expected to be very small throughout the experiments. Therefore, only the mean orientation errors ($E_{\text{q}}$ / $E_{\text{q}}^{*}$) are reported for the majority of this section for brevity.
@@ 5.3.1 Architecture & Backbone
Backbone Norm. Num. Param. Mem. [[[[MB]]]] Time [ms] $E_{\text{q}}$ / $E_{\text{q}}^{*}$ []
@@
This section begins with the study on OOD robustness and latency of various pose estimation architectures and NN backbones. First, Table 2 tabulates the latency and pose accuracy of different pose estimation architectures with various backbone NNs of different sizes. These backbones cover a wide variety of NN architectures, including conventional CNNs with BN layers (e.g., EfficientNet [20], EfficientNetV2 [123], MobileNetV3 [119], HRNet [101]); NFNet [125] which has no separate NORM layers but performs weight standardization of the convolution weights [127]; Vision Transformers (ViT) [22] with self-attention modules [128]; architectures purely based on Multi-Layer Perceptrons (MLP) (e.g., ResMLP [126]); and hybrid architectures which both include convolution operations and self-attention layers for image patches as hinted by the use of both BN/LN layers (e.g., HRFormer [120], EfficientFormerV2 [121], MobileViTv2 [122].
In order to facilitate readability, Table 2 sorts different architectures according to their total number of parameters in ascending order and grouped into similar sizes. Overall, larger NNs tend to result in better OOD robustness as measured by the performance on the SPEED+ lightbox and sunlamp domains. Then, by investigating the models within each size group, it becomes immediately clear that the EfficientDet (EfficientNet) and EfficientDet (EfficientNetV2) consistently outperform all other models across different size groups in terms of robustness. They are closely followed by EfficientDet (NFNet) and ViTPose (ViT).
Figure 6: Visualization of performance of different backbones tabulated in Table 2. Rotation error ($E_{\text{q}}^{*}$) is for the sunlamp domain.
Figure 6 condenses the information in Table 2 by plotting the mean orientation error on the sunlamp domain and the inference latency against the number of parameters. Different families of backbones are plotted with different colors and marker shapes. First, the $E_{\text{q}}^{*}$ plot in Fig. 6 (left) shows that EfficientDet (EfficientNet) and EfficientDet (EfficientNetV2) consistently outperform all other models in terms of OOD robustness. While EfficientDet (NFNet) models are closely behind them, there is no smaller variant of NFNet pre-trained on IN-1K that is publicly available. On the other hand, the latency plot in Fig. 6 (right) shows that ViTPose (ViT) and ViTPose (ResMLP) are clear winners. Both architectures consist solely of MLP layers, enabling a faster throughput compared to those with convolution operations. Unfortunately, ResMLP cannot perform well across domain gaps, which is no surprise since it is a pure MLP-based NN with no inductive bias of convolution operations or global spatial knowledge via self-attention layers. However, ViTPose (ViT) models closely follow EfficientDet (EfficientNet) in terms of OOD robustness.
Based on the OOD robustness and latency trade-off in Fig. 6, only EfficientDet (EfficientNet), EfficientDet (EfficientNetV2) and ViTPose (ViT) are considered for further analyses.
@@ 5.3.2 Latency on NVIDIA Jetson Nano
Table 3: Peak training memory and latency of select models from Table 2 are measured on an onboard GPU of NVIDIA Jetson Nano 4GB with PyTorch C++ API for batch size 1 and FP32 precision. The boldface denotes the best performance within each group.
Backbone Norm. Num. Param. Mem. [[[[MB]]]] Latency [ms] $E_{\text{q}}^{*}$ []
@@
The latency measured on the Jetson Nano reported in Table 3 reaffirms the observation made in Table 2 and Fig. 6. For the same number of parameters, EfficientDet models dominate on the OOD robustness front compared to the ViTPose models. However, ViTPose (ViT) models are consistently faster for inference regardless of the size. In fact, the largest ViT backbone considered in this study—ViT-M/16 with nearly 40M parameters—runs nearly 2.5 times faster than the smallest variant of the EfficientNet backbone that is 1/10 in size, and its training latency is on par with that of EfficientNet-B1 which only has 6.6M parameters. It also requires at most about 650 MB of GPU virtual memory for each training session with a single image input. Considering the 4GB of GPU memory available on the Jetson Nano and assuming that most of the GN&C-related operations run on CPUs, 650 MB or less is completely within the range of operation excluding the possibility of running other image-processing algorithms on the GPU.
Recall from the earlier discussion that the training is expected to run only occasionally whenever there has been a substantial change in the scenery and the target’s pose (section 3.1). Therefore, even the EfficientNet-B6 backbone could perform onboard online training asynchronously on a GPU given that there is enough power generated to support such a computationally heavy operation. In terms of the inference speed, all models considered in Table 3 can effectively run in real-time for 0.5 Hz GN&C updates of the CPOD reference mission [82] assuming that the pre- and post-processing of the NN inputs and outputs do not become the computational bottleneck.
@@ 5.3.3 Replacing BN Layers
Judging from the comparison on the NVIDIA Jetson Nano reported in Table 3, it appears that EfficientDet models are more OOD robust for the same number of parameters, but the larger ViTPose models are still more computationally efficient than small EfficientDet models, overcoming its comparative weakness in OOD robustness with increased size. There is one more reason that ViTPose is favored over EfficientDet, and that is due to the requirement that the onboard NNs must be batch-agnostic for online training (section 3.3). As evidenced in Table 2, most CNN backbones designed for ImageNet classification adopt BN layers, and even hybrid architectures always pair BN layers to each convolution operation. Then, is it possible to replace those BN layers with other batch-agnostic layers (e.g., LN, GN) when the overwhelming majority of publicly available CNN backbones pre-trained on ImageNet come with BN layers?
@@
Table 4 reports the orientation error of the EfficientDet (EfficientNet-B0) model (4.1M parameters) with different NORM layers during the IN-1K pre-training and the main training on SPEED+ synthetic images. The baseline is using BN layers for both sessions, which reports $E_{\text{q}}=6.63^{\circ}$ on lightbox and $E_{\text{q}}=10.77^{\circ}$ on sunlamp across 3 training sessions with different random seeds. When the BN layers of a pre-trained backbone is replaced with GN layers with group size 8, fully inheriting the trained affine parameters of the BN layers, the mean orientation error degrades by more than a factor of two. Training the network longer for 60 epochs instead of 30 does improve the overall performance, but it still does not reach the baseline level for all image domains. This was the conclusion of SPNv2 which also reported degradation in performance after replacing the BN with GN layers [18].
However, when the backbone is pre-trained on IN-1K with GN layers from the very beginning, the OOD performance shows consistent improvement on not only lightbox and sunlamp but also prisma25. The result indicates that it is possible to replace the BN layers in existing CNN architectures with GN or other batch-agnostic layers, but in order to preserve the OOD robustness, the backbone must be pre-trained on IN-1K or other large-scale datasets with those batch-agnostic layers from the very beginning. The problem is that ImageNet training is extremely expensive, both computationally and in terms of training time. For this study, EfficientNet-B0 was trained on 1M images of IN-1K loosely following the training Recipe B of Wightman et al. [129, Appendix F] for 350 epochs. Training this small NN on two NVIDIA RTX 4090 24GB GPUs took well over 100 hours. To train larger models would require even more computing power. If such heavy offline training can be afforded, one can use the EfficientNet-B0 backbone with GN layers that show OOD robustness comparable to the ViT-M/16 backbone with 40M parameters at the expense of increased inference time but with much less power consumption. However, if ImageNet pre-training cannot be afforded, it would be preferred to adopt ViT which is available for various ImageNet pre-trained models without BN layers.
@@ 5.3.4 Transfer Learning
The results so far indicate that ViTPose models are superior in terms of computational efficiency and batch-agnostic by nature, satisfying both Requirements #1 and #3 of section 3. However, they still fall short in OOD robustness when compared to EfficientDet models of comparable sizes. Therefore, it would be favorable to improve the OOD robustness of ViTPose models without substantial sacrifice of computational advantages. One potential remedy is to improve the transfer learning from the ImageNet pre-training. The ViT backbones of the ViTPose architecture so far are all pre-trained on IN-1K according to the DeiT III recipe [124]. This is an improved training method compared to DeiT [130] which is, in turn, an improved training method compared to the vanilla ViT [22]. In this section, the effect of ImageNet pre-training for ViT backbones is analyzed in detail by comparing the two different pre-training methods (DeiT vs. DeiT III) and datasets in Table 5. First, using random initialization without any pre-training on a large-scale dataset does not train well on SPEED+. For ViT-T/16 and ViT-S/16, using DeiT III leads to improved performance over DeiT when using only IN-1K for supervised pre-training. When the backbone is trained instead on a much larger ImageNet-22K (IN-22K) [34] which contains nearly 14M images and 22K class labels then fine-tuned on IN-1K, there is a marginal improvement in OOD robustness for both ViT-S/16 and ViT-M/16 models. The results overall indicate that better generalization of the backbone NN on a large-scale image classification dataset leads to a noticeable improvement in the downstream pose estimation task across the sim2real gap, though it seems the size of the pre-training dataset has little effect on the OOD robustness beyond the scale of IN-1K.
@@ 5.3.5 Input Resolution
Input Res. $\sigma$ [pix] Mem. [[[[MB]]]] Latency [ms] $E_{\text{q}}$ / $E_{\text{q}}^{*}$ []
@@
Another potential remedy to improve the OOD robustness of ViTPose models is to increase the input image resolution. The effect of input image resolution of ViTPose on the OOD robustness and latency on the Jetson Nano is reported in Table 6. Unsurprisingly, increasing the resolution of the input image cropped around the target leads to improved performance across the board but at the cost of increased peak training memory and latency. Specifically, doubling the input resolution from 224 $\times$ 224 to 448 $\times$ 448 nearly quadruples the training time for each input. This is expected, since doubling the input resolution but maintaining the same patch size (16 $\times$ 16) leads to double the length of patch tokens fed into the transformer blocks, and the memory footprint of the scaled dot-product operation of self-attention modules grows quadratically to the token length. Even though the training latency is tantamount to that of EfficientDet (EfficientNet-B6) operating on 224 $\times$ 224 inputs with nearly 40M parameters, the inference latency is still about a third and the training memory nearly half of EfficientDet (EfficientNet-B6) while performing better on all image domains. The result suggests that it could be beneficial to use ViT-based models with higher input resolutions as long as the training time does not become the bottleneck.

### v2

@@ 6.1 Datasets
Following SPEC2021 [21], this section performs experiments using the SPEED+ dataset [19, 119] which consists of 59,960 synthetic images split in 80:20 ratio into the training and validation sets. It further includes two HIL image domains acquired from the TRON facility and a mockup model of the Tango spacecraft: 6,740 lightbox and 2,791 sunlamp images which are reserved solely for testing. In this section, all NN models are trained on the synthetic training set and evaluated on the synthetic validation set, lightbox and sunlamp test sets, and 25 labeled flight images acquired in-space during RPO of the PRISMA mission (prisma25). Recall Fig.˜1 for visualization of a few HIL domain images.
@@ 6.2 Implementations
In the original EfficientDet implementation [106], the input image resolution and the size of the BiFPN and prediction layers differ depending on the EfficientNet scaling parameter $\phi$. In this section, for a fair comparison, the BiFPN layers are fixed to 4 convolution layers with 128 channels. The prediction heads for all pose estimation architectures consist of two convolution layers with 128 channels as well. Note that for most EfficientDet and HRNet implementations, the convolution operations in the prediction heads are depthwise separable convolutions [120] except ViTPose architectures that adopt full transposed 2D convolution operations with $\times$2 upsampling and 128 channels. The NORM and activation layers for BiFPN and prediction heads all follow those used in the respective backbone NN. If the backbone consists of multiple types of NORM layers including BN, the following BiFPN and prediction heads adopt only BN. The 2D keypoint locations are extracted from each heatmap as the location with maximum pixel intensity then refined via Unbiased Data Processing (UDP) [121]. The keypoint coordinates are used to solve for the 6D pose solution via EPnnP [87].
Given the ongoing interest in Artificial Intelligence (AI) for space applications, this work assumes that hardware support for edge computing to run NN onboard satellites will increase in the upcoming years (see section˜3.1.1). Therefore, this work adopts an NVIDIA Jetson Nano 4GB which contains a Quad-Core Arm® Cortex® A57 CPU and a 128-core NVIDIA Maxwell™ architecture GPU [27]. Nano is the most restrictive version of the commercially available Jetson family, which suits the limited processing power of onboard GPUs as evidenced in the NASA small spacecraft technology report [89]. Given its deprecated GPU architecture and CUDA language support, all NN training and inference on Jetson Nano are performed with full 32-bit Floating Point (FP32) precision using the C++ API of PyTorch 1.10.0.
6.2.1 Default Configuration
@@ 6.2.2 Metrics
[…]
@@
Note that in the overall pose estimation scenario (section˜4), the information on the relative position of the target is mostly included in the RoI detected from either the navigation filter or a separate object detection NN. Since the ground-truth RoIs are used for image pre-processing in this section, mean translation errors are expected to be very small throughout the experiments. Therefore, only the mean orientation errors ($E_{\text{q}}$ / $E_{\text{q}}^{*}$) are reported for the majority of this section for brevity.
@@ 6.3.1 Architecture & Backbone
This section begins with the study on OOD robustness and latency of various pose estimation architectures and NN backbones. First, Table˜2 tabulates the latency and pose accuracy of different pose estimation architectures with various backbone NNs of different sizes. These backbones cover a wide variety of NN architectures, including conventional CNNs with BN layers (e.g., EfficientNet [24], EfficientNetV2 [125], MobileNetV3 [126], HRNet [107]); NFNet [127] which has no separate NORM layers but performs weight standardization of the convolution weights [128]; Vision Transformers (ViT) [26, 129, 130] with self-attention modules [131]; architectures purely based on Multi-Layer Perceptrons (MLP) (e.g., ResMLP [132]); and hybrid architectures which both include convolution operations and self-attention layers for image patches as hinted by the use of both BN/LN layers (e.g., HRFormer [133], EfficientFormerV2 [134], MobileViTv2 [135].
@@
Backbone Norm. Num. Param. Mem. [[MB]] Time [ms] $E_{\text{q}}$ / $E_{\text{q}}^{*}$ []
@@
In order to facilitate readability, Table˜2 sorts different architectures according to their total number of parameters in ascending order and grouped into similar sizes. Overall, larger NNs tend to result in better OOD robustness as measured by the performance on the SPEED+ lightbox and sunlamp domains. Then, by investigating the models within each size group, it becomes immediately clear that the EfficientDet (EfficientNet) and EfficientDet (EfficientNetV2) consistently outperform all other models across different size groups in terms of robustness. They are closely followed by EfficientDet (NFNet) and ViTPose (ViT).
Figure 7: Visualization of performance of different backbones tabulated in Table˜2. Rotation error ($E_{\text{q}}^{*}$) is for the sunlamp domain.
Figure˜7 condenses the information in Table˜2 by plotting the mean orientation error on the sunlamp domain and the inference latency against the number of parameters. Different families of backbones are plotted with different colors and marker shapes. First, the $E_{\text{q}}^{*}$ plot in Fig.˜7 (left) shows that EfficientDet (EfficientNet) and EfficientDet (EfficientNetV2) consistently outperform all other models in terms of OOD robustness. While EfficientDet (NFNet) models are closely behind them, there is no smaller variant of NFNet pre-trained on IN-1K that is publicly available. On the other hand, the latency plot in Fig.˜7 (right) shows that ViTPose (ViT) and ViTPose (ResMLP) are clear winners. Both architectures consist solely of MLP layers, enabling a faster throughput compared to those with convolution operations. Unfortunately, ResMLP cannot perform well across domain gaps, which is no surprise since it is a pure MLP-based NN with no inductive bias of convolution operations or global spatial knowledge via self-attention layers. However, ViTPose (ViT) models closely follow EfficientDet (EfficientNet) in terms of OOD robustness.
Based on the OOD robustness and latency trade-off in Fig.˜7, only EfficientDet (EfficientNet), EfficientDet (EfficientNetV2) and ViTPose (ViT) are considered for further analyses.
@@ 6.3.2 Latency on NVIDIA Jetson Nano
Table 3: Peak training memory and latency of select models from Table˜2 are measured on an onboard GPU of NVIDIA Jetson Nano 4GB with PyTorch C++ API for batch size 1 and FP32 precision. The boldface denotes the best performance within each group.
Backbone Norm. Num. Param. Mem. [[MB]] Latency [ms] $E_{\text{q}}^{*}$ []
@@
The latency measured on the Jetson Nano reported in Table˜3 reaffirms the observation made in Table˜2 and Fig.˜7. For the same number of parameters, EfficientDet models dominate on the OOD robustness front compared to the ViTPose models. However, ViTPose (ViT) models are consistently faster for inference regardless of the size. In fact, the largest ViT backbone considered in this study—ViT-M/16 with nearly 40M parameters—runs nearly 2.5 times faster than the smallest variant of the EfficientNet backbone that is 1/10 in size, and its training latency is on par with that of EfficientNet-B1 which only has 6.6M parameters. It also requires at most about 650 MB of GPU memory for each training session with a single image input. Considering the 4GB of GPU memory available on the Jetson Nano and assuming that most of the GN&C-related operations run on CPUs, 650 MB or less is completely within the range of operation excluding the possibility of running other image-processing algorithms on the GPU.
Recall from the earlier discussion that the training is expected to run only occasionally whenever there has been a substantial change in the scenery and the target’s pose (section˜3.1). Therefore, even the EfficientNet-B6 backbone could perform onboard online training asynchronously on a GPU given that there is enough power generated to support such a computationally heavy operation. In terms of the inference speed, all models considered in Table˜3 can effectively run in real-time for 0.5 Hz GN&C updates of the CPOD reference mission [88] assuming that the pre- and post-processing of the NN inputs and outputs do not become the computational bottleneck.
@@ 6.3.3 Replacing BN Layers
Judging from the comparison on the NVIDIA Jetson Nano reported in Table˜3, it appears that EfficientDet models are more OOD robust for the same number of parameters, but the larger ViTPose models are still more computationally efficient than small EfficientDet models, overcoming its comparative weakness in OOD robustness with increased size. There is one more reason that ViTPose is favored over EfficientDet, and that is due to the requirement that the onboard NNs must be batch-agnostic for online training (section˜3.3). As evidenced in Table˜2, most CNN backbones designed for ImageNet classification adopt BN layers, and even hybrid architectures always pair BN layers to each convolution operation. Then, is it possible to replace those BN layers with other batch-agnostic layers (e.g., LN, GN) when the overwhelming majority of publicly available CNN backbones pre-trained on ImageNet come with BN layers?
@@
Table˜4 reports the orientation error of the EfficientDet (EfficientNet-B0) model (4.1M parameters) with different NORM layers during the IN-1K pre-training and the main training on SPEED+ synthetic images. Note that it reports the mean and standard deviations of average rotation errors across three separate training sessions with different random seeds. This is in order to demonstrate that the reported performances are not “lucky” results due to some favorable random seeds, but rather all training sessions lead to consistent minima of the training objective. The baseline is using BN layers for both sessions, which reports $E_{\text{q}}=6.63^{\circ}$ on lightbox and $E_{\text{q}}=10.77^{\circ}$ on sunlamp across 3 training sessions with different random seeds. When the BN layers of a pre-trained backbone is replaced with GN layers with group size 8, fully inheriting the trained affine parameters of the BN layers, the mean orientation error degrades by more than a factor of two. Training the network longer for 60 epochs instead of 30 does improve the overall performance, but it still does not reach the baseline level for all image domains. This was the conclusion of SPNv2 which also reported degradation in performance after replacing the BN with GN layers [22].
However, when the backbone is pre-trained on IN-1K with GN layers from the very beginning, the OOD performance shows consistent improvement on not only lightbox and sunlamp but also prisma25. The result indicates that it is possible to replace the BN layers in existing CNN architectures with GN or other batch-agnostic layers, but in order to preserve the OOD robustness, the backbone must be pre-trained on IN-1K or other large-scale datasets with those batch-agnostic layers from the very beginning. The problem is that ImageNet training is extremely expensive, both computationally and in terms of training time. For this study, EfficientNet-B0 was trained on 1M images of IN-1K loosely following the training Recipe B of Wightman et al. [136, Appendix F] for 350 epochs. Training this small NN on two NVIDIA RTX 4090 24GB GPUs took well over 100 hours. To train larger models would require even more computing power. If such heavy offline training can be afforded, one can use the EfficientNet-B0 backbone with GN layers that show OOD robustness comparable to the ViT-M/16 backbone with 40M parameters at the expense of increased inference time but with much less power consumption. However, if ImageNet pre-training cannot be afforded, it would be preferred to adopt ViT which is available for various ImageNet pre-trained models without BN layers.
@@ 6.3.4 Transfer Learning
The results so far indicate that ViTPose models are superior in terms of computational efficiency and are batch-agnostic by nature, satisfying both Requirements #1 and #3 of section˜3. However, they still fall short in OOD robustness when compared to EfficientDet models of comparable sizes. Therefore, it would be favorable to improve the OOD robustness of ViTPose models without substantial sacrifice of computational advantages. One potential remedy is to improve the transfer learning from the ImageNet pre-training. The ViT backbones of the ViTPose architecture so far are all pre-trained on IN-1K according to the DeiT III recipe [130]. This is an improved training method compared to DeiT [129] which is, in turn, an improved training method compared to the vanilla ViT [26]. In this section, the effect of ImageNet pre-training for ViT backbones is analyzed in detail by comparing the two different pre-training methods (DeiT vs. DeiT III) and datasets in Table˜5. First, using random initialization without any pre-training on a large-scale dataset does not train well on SPEED+. For ViT-T/16 and ViT-S/16, using DeiT III leads to improved performance over DeiT when using only IN-1K for supervised pre-training. When the backbone is trained instead on a much larger ImageNet-22K (IN-22K) [39] which contains nearly 14M images and 22K class labels then fine-tuned on IN-1K, there is a marginal improvement in OOD robustness for both ViT-S/16 and ViT-M/16 models. The results overall indicate that better generalization of the backbone NN on a large-scale image classification dataset leads to a noticeable improvement in the downstream pose estimation task across the sim2real gap, though it seems the size of the pre-training dataset has little effect on the OOD robustness beyond the scale of IN-1K.
@@ 6.3.5 Input Resolution
Input Res. $\sigma$ [pix] Mem. [[MB]] Latency [ms] $E_{\text{q}}$ / $E_{\text{q}}^{*}$ []
@@
Another potential remedy to improve the OOD robustness of ViTPose models is to increase the input image resolution. The effect of input image resolution of ViTPose on the OOD robustness and latency on the Jetson Nano is reported in Table˜6. Unsurprisingly, increasing the resolution of the input image cropped around the target leads to improved performance across the board but at the cost of increased peak training memory and latency. Specifically, doubling the input resolution from 224 $\times$ 224 to 448 $\times$ 448 nearly quadruples the training time for each input. This is expected, since doubling the input resolution but maintaining the same patch size (16 $\times$ 16) leads to double the length of patch tokens fed into the transformer blocks, and the memory footprint of the scaled dot-product operation of self-attention modules grows quadratically to the token length. Even though the training latency is tantamount to that of EfficientDet (EfficientNet-B6) operating on 224 $\times$ 224 inputs with nearly 40M parameters, the inference latency is still about a third and the training memory nearly half of EfficientDet (EfficientNet-B6) while performing better on all image domains. The result suggests that it could be beneficial to use ViT-based models with higher input resolutions as long as the training time does not become the bottleneck.